Find the Cartesian equation of the locus of the point P representing the complex number z. Sketch the locus of P in each case.
1
Expert's answer
2020-05-27T18:50:02-0400

a)

"2|z+1|=|z-2|"

"z=x+iy"


"4((x+1)^2+y^2)=(x-2)^2+y^2""4x^2+8x+4+4y^2=x^2-4x+4+y^2""3x^2+12x+3y^2=0""x^2+4x+4-4+y^2=0""(x+2)^2+y^2=4"

The equation of the circle with center "(-2,0)" and radius "2."


b)


"|{z+i \over z-5-2i}|=1""|z+i|=|z-5-2i|,z\not=5+2i"

"z=x+iy"


"x^2+(y+1)^2=(x-5)^2+(y-2)^2""x^2+y^2+2y+1=x^2-10x+25+y^2-4y+4""6y=-10x+28"

"y=-{5 \over 3}x+{14 \over 3}"


c)


"Im(z+{9\over z})=0"

"z=x+iy, z\not=0"


"x+iy+{9\over x+iy}=x+iy+{9\over x^2+y^2}(x-iy)"

"y(x^2+y^2)-9y=0"

"y=0 \ or \ x^2+y^2=9"

Real axis or the circle with center "(0,0)" and radius "3" with the exception of point "(0,0)."